OpenStudy (anonymous):

why would the following solution space have no basis: 2x_1+x_2+3x_3=0 x_1 + 0 +5x_3=0 0 + x_2+x_3=0 I know it reduces to 1 0 0 0 1 0 0 0 1 but why does that mean it has no basis? I would think that a 3x3 identity matrix of 3 vectors i, j ,k can span R^3

OpenStudy (anonymous):

hey can you help me out later?

ganeshie8 (ganeshie8):

it is a full rank matrix so Ax=0 will be having only the trivial solution : (0, 0, 0)

ganeshie8 (ganeshie8):

what do you know about the nullspace of a matrix ?

OpenStudy (anonymous):

Im not too sure what youre trying to ask... I would say that a nullspace is a homogeneous system?

ganeshie8 (ganeshie8):

nullspace is one of the four fundamenal spaces of a matrix, it is a set of all the solutions of Ax=0

OpenStudy (anonymous):

Ah I see so the solutions to the Ax=0 matrix would be a nullspace. My book doesn't talk about nullspace, more like solution space / basis ( it interchanges them).

ganeshie8 (ganeshie8):

"solution space" and "basis" are two different things

ganeshie8 (ganeshie8):

basis of a vector space is just a set of linearly independent vectors whose linear combinations span the space for example, the column vectors in 3x3 identity matric form a "basis" for the vector space R^n

OpenStudy (anonymous):

Nvm what I just said, it asked me to find the basis of a solution space*

OpenStudy (anonymous):

Then, if I solve for a solution space and it gives a 3x3 identity matrix, shouldn't it give me the basis of the solution space too?

ganeshie8 (ganeshie8):

those vectors in 3x3 identity matrix form a basis of "column space"

ganeshie8 (ganeshie8):

full rank just tells you that you can reach any vector in R^3 by taking linear combinations of the column vectors : Ax = b has an unique solution for all vectors b

ganeshie8 (ganeshie8):

the homogeneous system Ax = 0 will have a trivial solution only so the dimension of nullspace would be 0 and we say there is "no basis"

OpenStudy (anonymous):

I don't see why you're talking about Ax=b. I mean a \solution space will always be homogeneous no?

ganeshie8 (ganeshie8):

\[\begin{bmatrix}2&1&3\\1&0&5\\0&1&1\end{bmatrix}\begin{pmatrix}x_1\\x_2\\x_3\end{pmatrix} = \begin{pmatrix}0\\0\\0\end{pmatrix}\] equivalantly, you could also say that the column vectors are linearly independent, so the above system will only have a trivial solution : x1 = x2 = x3 = 0

ganeshie8 (ganeshie8):

Notice, the solution space of Ax = 0 will not span R^n. it only contains the zero vector.

OpenStudy (anonymous):

OH I see, if it is linearly independent in means the only coordinate vector would be (0,0,0) so it would only give a dimension of 0 which doesn't have a basis